# GUIDE: Towards Scalable Advising for Research Ideas

### v1

@@ Abstract
Yaowenqi Liu, BingXu Meng, Rui Pan, Jerry Huang, Tong Zhang University of Illinois Urbana-Champaign {yl140, bingxum2, ruip4, jerry8, tozhang}@illinois.edu
@@ 3.2 Data Collection & Generation
An essential step in our preprocessing pipeline is content compression through summary generation. To this end, we used gpt-4.1-nano, a cost-effective yet high-performing model from OpenAI (Achiam et al., 2023), to generate structured section-wise summaries for each paper (e.g., Introduction, Related Work, Methodology, Experiments). This summarization reduced the input length of each paper by approximately 16×16$\times$16 ×, allowing us to incorporate substantially more context within the RAG input window, mitigating the token limit bottleneck and improving retrieval efficiency in downstream tasks.
@@ 3.4 Reward Ranked Fine-Tuning
LLMs often suffer from implicit biases, leading to suboptimal or skewed outputs. In our experiments, we observed that off-the-shelf models tend to offer only positive and overly general praise and rarely provide neutral or critical feedback. This phenomenon highlights the need for better alignment with human evaluative standards. To improve the alignment of our models, we use Reward-rAnked Fine-Tuning (RAFT; Dong et al., 2023), an iterative fine-tuning algorithm with rejection sampling. Details of the pipeline are in illustrated in Figure 3.
@@
To empower general-purpose small LLMs to learn advising-centered reasoning and output formats, we adopt a warm-up phase at the start of training. Rubrics-prompted DeepSeek-R1 (Guo et al., 2025) is employed to generate evaluations for a randomly sampled subset of ICLR 2024 papers, producing 4,000 high-quality idea–evaluation pairs. These examples are used to perform an initial round of supervised fine-tuning (SFT) of Qwen2.5-7B-Instruct.
@@
To further reduce learning difficulty, we introduce an additional text-similarity reward $R^{\mathrm{text}}_{i}$ by measuring the ROUGE score (Lin, 2004) between the generated advice and the concatenation of all reference human reviews. The overall reward is then given by
@@
where $\lambda\in[0,1]$ balances the two objectives. We select the candidate advice with the highest $R_{i}$ at each iteration for supervised fine-tuning.
@@
After computing the combined reward, the advice $a_{i}^{*}$ with highest reward among $K$ candidates are selected, which forms a supervised fine-tuning set $\mathcal{S}=\{(x_{i},a_{i}^{*})\}_{i}$. Here $x_{i}$ denote the retrieval augmented input for hypothesis $i$, i.e. the concatenation of the paper’s abstract, claimed contributions, method description, experimental setup, and the retrieved summaries from Sec 3.3. Qwen2.5-7B-Instruct is fine-tuned on $\mathcal{S}$. By iterating this generate–select–fine-tuning cycle, the model progressively learns to produce advices that maximize alignment with human judgments and textual fidelity.
@@ 4.1 Experiment Setting
We compare GUIDE with baselines using various large general-purpose LLMs, including GPT-4o-mini, QwQ-32B, and DeepSeek-R1, all equipped with retrieval-augmented generation, as described in Sec. 3.3. For all LLMs, the decoding hyperparameters are set to temperature $=0.6$ and top-$p$ with $p=0.95$. Both GUIDE-7B and baselines will receive the input hypothesis’s abstract, claimed contribution, method description, and experimental setup, along with the ten most relevant literature sections from our database.
@@ 4.4 Rubrics Guided Prompting
In this experiment, we fix the retrieved contents to be the same, all with 10 related abstracts, 10 contributions, 10 method summaries, and 10 experimental setups. The only difference across variants is the system prompt, which directs the LLM to emphasize a specific rubric (e.g., significance, novelty, soundness) and output the corresponding aspect-focused evaluation. Since QwQ-32B exhibits relatively weaker instruction-following capabilities under prompting, it is replaced with another widely adopted Gemini-flash-2.0 to ensure robust adherence to our system prompts.
@@ Results
As shown in Table 4, we observe that the significance rubric yields the most pronounced improvement in evaluation quality. To demonstrate the effectiveness of this rubric-guided approach, a concrete example is presented to illustrate the model’s assessment with and without significance rubrics applied to an ICLR2025 Oral hypothesis, as shown in Table 5. For more detailed rubrics based prompting, please refer to Appendix C.4.
@@ 4.5 Uncertainty Analysis
A high entropy indicates that the classifier’s probability mass is spread across many rating classes, whereas a low entropy reflects a focused, high-confidence prediction.
To assess the impact of prediction confidence on evaluation accuracy, we rank all hypotheses by ascending entropy and select three subsets corresponding to the lowest-entropy: 10%, 20%, and 30% of papers. Within each subset, we recompute the Top-30% Precision metric, measuring the proportion of truly accepted papers among the top 30% of model-ranked hypotheses.
@@ Results
Fig. 4 reveals how model confidence modulates evaluation reliability: as the confidence and ranking threshold tightens, we generally observe higher precision, indicating that low-entropy predictions are more trustworthy indicators of acceptance. Notably, within the top 10% confidence subset, the Top-30% Precision reaches the high accuracy of 93.3%, suggesting that even for hypotheses not yet fully formalized into manuscripts, meeting both high-confidence and high-ranking criteria is a strong indicator of the final acceptance. In automated hypothesis generation pipelines, where large batches of hypotheses can be generated at low cost, leveraging uncertainty enables the automatic selection of high-quality hypotheses, underscoring the potential for accelerating research discovery.
@@ 5 Discussion
Although our model and system are specifically tailored for academic hypotheses advising rather than full-text academic paper review, in direct comparisons they outperform existing general-purpose LLM-based baselines and achieve precision levels that closely approximate those of human reviewers.
@@
For a strong and well-known baseline, we employ the review agent component of AI Scientist (Lu et al., 2024). It processes the entire paper through multiple rounds of reflection and aggregates outputs via an ensemble of model passes to produce its final evaluation. By comparing against this full-text, multi-stage baseline, we demonstrate the advantages of our hypotheses-centric advisor. Since GPT-4o-mini tends to reject all papers under AI Scientist’s default settings, we replaced it with GPT-4.1-nano to ensure stable results.
@@ Results
Table 6: Performance comparison on the ICLR 2025 test set, where Accuracy (or Precision) =#$=\frac{\#\text{Correct Predicted Acceptances}}{\#\text{Predicted Acceptances}}$ represents the proportion of correct accept decisions over all accept decisions made by the system, and F1 is the harmonic mean of Precision and Recall, with the Recall =#$=\frac{\#\text{Correct Predicted Acceptances}}{\#\text{Actual Acceptances}}$ being the proportion of correct accept decisions made by the system over all the actual accepted papers.
Table 5 reveals that our idea-centric advisor achieves similar performance as the human baseline in terms of acceptance rate. Moreover, our system outperforms the AI Scientist’s review agent, which exhibits strong decision biases: GPT-4.1-nano accepts 17.1% of papers, DeepSeek-R1 accepts 85.4% of papers, and QwQ-32B accepts 69.2%, all far away from the true 31.7% rate. These results underscore the advantage of using a ranking-based evaluation standard, which more faithfully reflects selective thresholds and yields more balanced, reliable assessments.
@@ C.3 Rating Distribution
Table C.3 shows that human reviewers give the lowest average score (5.13) but with the greatest spread in their judgments (variance = 1.46). All LLM-based evaluators assign higher mean ratings, yet their variances (0.10–0.26) are much lower, indicating they rate more consistently but also more leniently. Among them, GUIDE-7B comes closest to human behavior, with a mean of 5.81 and a variance of 0.22.
@@ E.1 Training
  • Generation: For each iteration, we use vLLM to sample 1,000 ICLR 2024 papers and generate $K=16$ candidate evaluations per hypothesis with temperature $0.7$, top-$p$ $=0.8$, and repetition penalty $=1.05$.
@@
    The text-similarity reward $R^{\mathrm{text}}$ is the sum of ROUGE-1, ROUGE-2, and ROUGE-L, each weighted by 0.1 (total weight 0.3).
  • Supervised Fine-Tuning: We keep fine-tuning our warmed-up model via LoRA with rank $r=64$ and alpha $=64$, using learning rate $1\times 10^{-5}$, batch size $16$, for 2 epochs, and a cosine learning-rate schedule. We set the context window equal to 15k tokens. Training runs on 2 NVIDIA A100 40 GB GPUs with DeepSpeed’s ZeRO 3 optimizations and CPU offload enabled.

### v2

@@ Abstract
Yaowenqi Liu, Bingxu Meng, Rui Pan, Yuxing Liu, Jerry Huang, Jiaxuan You, Tong Zhang University of Illinois Urbana-Champaign {yl140, bingxum2, ruip4, yuxing6, jerry8, jiaxuan, tozhang}@illinois.edu
@@ 3.2 Data Collection & Generation
An essential step in our preprocessing pipeline is content compression through summary generation. To this end, we used gpt-4.1-nano, a cost-effective yet high-performing model from OpenAI (Achiam et al., 2023), to generate structured section-wise summaries for each paper (e.g., Introduction, Related Work, Methodology, Experiments). This summarization reduced the input length of each paper by approximately $16\times$, allowing us to incorporate substantially more context within the RAG input window, mitigating the token limit bottleneck and improving retrieval efficiency in downstream tasks.
@@ 3.4 Reward Ranked Fine-Tuning
LLMs often suffer from implicit biases, leading to suboptimal or skewed outputs. In our experiments, we observed that off‐the‐shelf models tend to offer only positive and overly general praise and rarely provide neutral or critical feedback. This phenomenon highlights the need for better alignment with human evaluative standards. To improve the alignment of our models, we use Reward‐rAnked Fine‐Tuning (RAFT; Dong et al., 2023), an iterative fine-tuning algorithm with rejection sampling. Details of the pipeline are in illustrated in Figure 3.
@@
To empower general-purpose small LLMs to learn advising-centered reasoning and output formats, we adopt a warm-up phase at the start of training. Rubrics-prompted DeepSeek‐R1 (Guo et al., 2025) is employed to generate evaluations for a randomly sampled subset of ICLR 2024 papers, producing 4,000 high‐quality idea–evaluation pairs. These examples are used to perform an initial round of supervised fine‐tuning (SFT) of Qwen2.5‐7B‐Instruct.
@@
To further reduce learning difficulty, we introduce an additional text‐similarity reward $R^{\mathrm{text}}_{i}$ by measuring the ROUGE score (Lin, 2004) between the generated advice and the concatenation of all reference human reviews. The overall reward is then given by
@@
where $\lambda\in[0,1]$ balances the two objectives. We select the candidate advice with the highest $R_{i}$ at each iteration for supervised fine‐tuning.
@@
After computing the combined reward, the advice $a_{i}^{*}$ with highest reward among $K$ candidates are selected, which forms a supervised fine‐tuning set $\mathcal{S}=\{(x_{i},a_{i}^{*})\}_{i}$. Here $x_{i}$ denote the retrieval augmented input for hypothesis $i$, i.e. the concatenation of the paper’s abstract, claimed contributions, method description, experimental setup, and the retrieved summaries from Sec 3.3. Qwen2.5‐7B‐Instruct is fine‐tuned on $\mathcal{S}$. By iterating this generate–select–fine‐tuning cycle, the model progressively learns to produce advices that maximize alignment with human judgments and textual fidelity.
@@ 4.1 Experiment Setting
We compare GUIDE with baselines using various large general-purpose LLMs, including GPT-4o-mini, QwQ-32B, and DeepSeek-R1, all equipped with retrieval-augmented generation, as described in Sec. 3.3. For all LLMs, the decoding hyperparameters are set to temperature =0.6=0.6 and top-$p$ with $p=0.95$. Both GUIDE-7B and baselines will receive the input hypothesis’s abstract, claimed contribution, method description, and experimental setup, along with the ten most relevant literature sections from our database.
@@ 4.4 Rubrics Guided Prompting
In this experiment, we fix the retrieved contents to be the same, all with 10 related abstracts, 10 contributions, 10 method summaries, and 10 experimental setups. The only difference across variants is the system prompt, which directs the LLM to emphasize a specific rubric (e.g., significance, novelty, soundness) and output the corresponding aspect‐focused evaluation. Since QwQ-32B exhibits relatively weaker instruction-following capabilities under prompting, it is replaced with another widely adopted Gemini-flash-2.0 to ensure robust adherence to our system prompts.
@@ Results
As shown in Table 4, we observe that the significance rubric yields the most pronounced improvement in evaluation quality. To demonstrate the effectiveness of this rubric‐guided approach, a concrete example is presented to illustrate the model’s assessment with and without significance rubrics applied to an ICLR2025 Oral hypothesis, as shown in Table 5. For more detailed rubrics based prompting, please refer to Appendix C.4.
@@ 4.5 Uncertainty Analysis
A high entropy indicates that the classifier’s probability mass is spread across many rating classes, whereas a low entropy reflects a focused, high‐confidence prediction.
To assess the impact of prediction confidence on evaluation accuracy, we rank all hypotheses by ascending entropy and select three subsets corresponding to the lowest‐entropy: 10%, 20%, and 30% of papers. Within each subset, we recompute the Top‐30% Precision metric, measuring the proportion of truly accepted papers among the top 30% of model‐ranked hypotheses.
@@ Results
Fig. 4 reveals how model confidence modulates evaluation reliability: as the confidence and ranking threshold tightens, we generally observe higher precision, indicating that low‐entropy predictions are more trustworthy indicators of acceptance. Notably, within the top 10% confidence subset, the Top-30% Precision reaches the high accuracy of 93.3%, suggesting that even for hypotheses not yet fully formalized into manuscripts, meeting both high‐confidence and high‐ranking criteria is a strong indicator of the final acceptance. In automated hypothesis generation pipelines, where large batches of hypotheses can be generated at low cost, leveraging uncertainty enables the automatic selection of high‐quality hypotheses, underscoring the potential for accelerating research discovery.
@@ 5 Discussion
Although our model and system are specifically tailored for academic hypotheses advising rather than full‐text academic paper review, in direct comparisons they outperform existing general-purpose LLM‐based baselines and achieve precision levels that closely approximate those of human reviewers.
@@
For a strong and well‐known baseline, we employ the review agent component of AI Scientist (Lu et al., 2024). It processes the entire paper through multiple rounds of reflection and aggregates outputs via an ensemble of model passes to produce its final evaluation. By comparing against this full‐text, multi‐stage baseline, we demonstrate the advantages of our hypotheses‐centric advisor. Since GPT-4o-mini tends to reject all papers under AI Scientist’s default settings, we replaced it with GPT-4.1-nano to ensure stable results.
@@ Results
Table 6: Performance comparison on the ICLR 2025 test set, where Accuracy (or Precision) $\frac{\#\text{Correct Predicted Acceptances}}{\#\text{Predicted Acceptances}}$ represents the proportion of correct accept decisions over all accept decisions made by the system, and F1 is the harmonic mean of Precision and Recall, with the Recall $\frac{\#\text{Correct Predicted Acceptances}}{\#\text{Actual Acceptances}}$ being the proportion of correct accept decisions made by the system over all the actual accepted papers.
Table 6 reveals that our idea‐centric advisor achieves similar performance as the human baseline in terms of acceptance rate. Moreover, our system outperforms the AI Scientist’s review agent, which exhibits strong decision biases: GPT-4.1-nano accepts 17.1% of papers, DeepSeek-R1 accepts 85.4% of papers, and QwQ-32B accepts 69.2%, all far away from the true 31.7% rate. These results underscore the advantage of using a ranking-based evaluation standard, which more faithfully reflects selective thresholds and yields more balanced, reliable assessments.
@@ C.3 Rating Distribution
Table 13 shows that human reviewers give the lowest average score (5.13) but with the greatest spread in their judgments (variance = 1.46). All LLM-based evaluators assign higher mean ratings, yet their variances (0.10–0.26) are much lower, indicating they rate more consistently but also more leniently. Among them, GUIDE-7B comes closest to human behavior, with a mean of 5.81 and a variance of 0.22.
@@ E.1 Training
  • Generation: For each iteration, we use vLLM to sample 1,000 ICLR 2024 papers and generate $K=16$ candidate evaluations per hypothesis with temperature $0.7$, top-$p$ =0.8=0.8, and repetition penalty =1.05=1.05.
@@
    The text‐similarity reward $R^{\mathrm{text}}$ is the sum of ROUGE‐1, ROUGE‐2, and ROUGE‐L, each weighted by 0.1 (total weight 0.3).
  • Supervised Fine-Tuning: We keep fine-tuning our warmed-up model via LoRA with rank $r=64$ and alpha =64=64, using learning rate $1\times 10^{-5}$, batch size $16$, for 2 epochs, and a cosine learning-rate schedule. We set the context window equal to 15k tokens. Training runs on 2 NVIDIA A100 40 GB GPUs with DeepSpeed’s ZeRO 3 optimizations and CPU offload enabled.